A galactic police officer's hovercraft is parked a
Answer #1

Initial velocity of hovercraft = vih= 0 m/s ……(It’s correct)

Initial position of speederbike = x0s = 68*2 = 136 m

a) xfh = xfs

x0h+vih*t +1/2*ah*t^2 = x0s+vis*t

0+0*t -1/2*4*t^2 = 136+86*t   => t= 35.9s

xfh=xfs = 2577.2m

b) The maximum distance = 68.0*2.0 = 136 m